OpenStudy (anonymous):

I am new to this, can someone help? The widthof a rectangle is 4cm less than the length. the perimeter is at most 48 cm. What are the restrictions on the dimensions of the rectangle?

OpenStudy (anonymous):

L = w + 4 P = 2 ( w + L ) = 48 -> w + w + 4 = 24 => 2w = 20 Can you continue?